Two jets leave an air base at the same time and travel in opposite directions. One jet travels 92 slower than the other. If the two jets are 10,602 apart after 9 hours, what is the rate of each jet?
Found 2 solutions by ewatrrr, rfadrogane:
Answer by ewatrrr(10682)


Hi
Let x and (x-92)represent the rates each of the jets are traveling
Question states*** D= r*t
9x + 9(x-92) = 10,602mi
Solving for x
18x = 10602+ 825
18x = 11430
x = 635mph. The other jet traveling at 543mph. (635-92)
CHECKING our Answer***
9hr*635mph + 9hr*543mph = 10,602mi

Answer by rfadrogane(214)

Two jets leave an air base at the same time and travel in opposite directions. One jet travels 92 slower than the other. If the two jets are 10,602 apart after 9 hours, what is the rate of each jet?
Sol'n:
note: your lacking of units...
St=S1+S2
10,602=9v+9(v-92)
10,602=18v-828
18v=10,602+828
18v=11,430
v=635
v-92=543

thus the rate are 635 & 543 ----answer
^_^